online updates

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Sun Aug 26 2001 - 20:41:29 EDT

  • Next message: Apache: "Shift report 20010827 16:00-24:00"

    A number of changes in the daq config & control scripts  etc has been made over the last couple of dates. This is a brief summary
    
    -software
    a) A rawdata record of H1Cal slats has been added to the mapped data. It gives tdc+adc for the two cal slats behind C1/H1
       It still has to be implemented in the BrRawDatainput.
    b) A record to check the trigger mask in fb3 was added
    c) A new monitorKeithley script was written by the Bergen students and implemented.
    d) A new script to get lists of runinfo was added (Bergen students)
    e) The fb4 'list'record was modified to get rid of front end errors.
    - hardware
    a) To chekc out the zdc performance and help in deciding on a better configuration the hardware for Yellow beam was modified today
       the ZDC1-3 input were split before the linear fan i/o with a signal directly to adc , and the other via a capacitive coupling. 
       It should be checked what effect this have on the energy signal of (high gain left) and other perf.
       The coinc rates seems form the short run at noon today  not to be efected.
     
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Sun Aug 26 2001 - 20:33:25 EDT